OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4)

OpenStudy (anonymous):

2 = 7m + b 4 = 5m + b then m = -1 b = 9 so the equation should be: y = -x + 9